# Towards the Pedagogical Steering of Large Language Models for Tutoring: A Case Study with Modeling Productive Failure

### v1

@@ A.4 Intent Dependent Steering
type grade time name reference exercise solution
Type between invention, ill-structured, and well-structured Recommended school grade Recommended time to complete the exercise Name to quickly identify the exercise Paper from which the exercise was curated Exercise statement Solution if provided in paper
Table 5: Description of the fields in the dataset.
A.5 Learning Sciences Rationales of our Productive Failure Modelling
The Intent Selection graph presented in Figure 4 has been created from expert knowledge in Productive Failure. How to read this graph is explained in Appendix A.3. We detail in this paragraph some of the principles we used to create it.
In Productive Failure, the goal is to make the student explore the solution space of an open-ended problem as much as possible, even if they fail to produce a complete solution. This idea gave its name to Productive Failure: there is, in this case, a hidden efficacy to the student’s failure (Kapur 2008). To this aim, in Figure 4 any mistake (features a, b or c in Figure 3) leads to a Guide Self Correction tutoring intent. The goal is to make the student spot and correct their mistake by themself. If the error persists, we use the intent Correct that gives a hint to the student about their mistake, but no intent seeks to directly correct it. If the mistake is computational only, we use the intent Offload that prompts the LLM to perform the computation. This allows the student to focus on exploring the solution space rather than wasting time on computations. For the same reason, if the student asks for a basic theorem or definition they forgot (feature h), we use the intent State to answer their question and help them focus on the reasoning.
A good tutor should intervene very indirectly (Merrill et al. 1992): a good principle to follow as a tutor is to give as little scaffolding as one can get away with. The intent used whenever a student manages to correct their mistake (arrow labeled f leaving the Guide Self-Correction node in Figure 4) or when the student suggests an intuitive or incomplete solution (state feature d) is the Seek Strategy intent. This encourages the student to pursue their line of thought without heavier scaffolding.
One of the challenges of Productive Failure is to create a classroom culture that allows for failure and promotes psychological safety. Indeed, as students deal with open-ended problems before being given the tools to solve them, the experience can be stressful. In traditional classroom settings, the focus is directed towards getting the right solution and students are often evaluated on this with a grade. One of the most crucial design features of a Productive Failure problem session is to shift this goal from solving the problem to generating multiple RSMs regardless of their correctness (Kapur and Bielaczyc 2012). As explained in Kapur and Bielaczyc (2012), “students are used to seeking assistance from their teachers so much so that they do so even before sufficiently trying to solve problems themselves. At the same time, teachers are just as used to providing assistance when it is asked for so much so that often opportunities for students to generate and explore RSMs are missed”. For this reason, in our modelling, if the student explicitly asks for help the intent Prompt Intuition/Hypothesis is triggered. This intent encourages the student to start with a guess or an intuition. This pushes students to try with no fear of being incorrect. Once the student is successful in providing an intuitive solution, or if they confidently propose an incomplete one (feature e), then the Elicit Articulation intent is used to encourage an RSM creation by formulating their idea in concrete math. To mitigate stress, that can hinder the student’s learning (Lepper et al. 1993), we trigger a Hint intent to ease the student’s progress if they show signs of lack of motivation or self-confidence.
A good Productive Failure problem allows for multiple solutions (and often a preferred one, called canonical solution). Then, if the student manages to propose a non-canonical solution to the problem (feature j), they should be encouraged to continue exploring. The Identify Limits intent is triggered to ask the student to identify the limitations of their solution, encouraging them to develop a new one.
If the student reaches the canonical solution (feature k), we use the intent (Self-)Reflect. It has been proven that it is beneficial from a learning point of view to engage in self-reflection (i.e. looking back critically at the work done) after a problem session (Katz, O’Donnell, and Kay 2000). Following this turn, the tutor is made ready to end the conversation through the State Greetings/Farewell intent.
A.6 Dataset
We collected 20 problems from the Productive Failure literature  (Kapur 2008; Kapur and Bielaczyc 2012; Kapur 2010; Kerrigan, Weber, and Chinn 2021; Kerrigan 2018). We present the columns of the dataset in Table 5. We curated three types of problems: Invention problems ask the student to invent a new way to solve a problem. They typically admit several correct but suboptimal answers, as long as the student can justify them. Students are not expected to have previous exposure to the tools canonically used to solve the exercise. The problem Consistency is a good example of invention problem. Ill-structured problems are open-ended problems that contain some non-relevant information and a convoluted story. These problems often admit a single solution, but there can be several ways to reach it. The problem country is an example of such a problem. Well-structured problems are more direct versions of the ill-structured problems. They resemble classical exercises that students usually solve in a mathematics class. More information about the different types of exercise in Productive Failure and their role can be found in (Kapur 2008) and (Kapur and Bielaczyc 2012).
A.7 Test Problems
Figures 6 and 7 display the two test problems we used in our field test. They are also included in the dataset we release.
A.8 PF Score Details
We present in the following paragraphs the precise reasoning sub-steps used for the Productive Failure Score grading as explained in Section 4.4.
Show your work and justify why you think your method is correct.
Solution:
Initial quotas:
- A: 1,646,000 / 50,000 = 32.92 → 32 seats
- B: 6,936,000 / 50,000 = 138.72 → 138 seats
- C: 154,000 / 50,000 = 3.08 → 3 seats
- D: 2,091,000 / 50,000 = 41.82 → 41 seats
- E: 685,000 / 50,000 = 13.70 → 13 seats
- F: 988,000 / 50,000 = 19.76 → 19 seats
Total initial seats assigned = 32 + 138 + 3 + 41 + 13 + 19 = 246
Seats left to distribute = 250 - 246 = 4
Distribute the surplus seats based on largest remainders:
Remainders (from the divisions above):
- A: 0.92
- B: 0.72
- C: 0.08
- D: 0.82
- E: 0.70
- F: 0.76
Figure 6: Problem Country and its solution.
The organizers of the Premier League Federation have to decide which one of the three players Mike Arwen, Dave Backhand and Ivan Right - should receive the ”The Most Consistent Player for the Past 5 Years” award. Table 1 shows the number of goals that each striker scored between 2019 and 2023.
(1) Design as many different measures of consistency as you can.
Year Mike Arwen Dave Backhand Ivan Right
2007 13 12 14
2008 12 14 10
2009 15 16 18
2010 17 15 18
2011 13 13 15
Solution:
First, compute the mean:
Mean number of goals for Mike: 14
Mean number of goals for Dave: 14
Mean number of goals for Ivan: 15
Sum squared deviation for Mike: 16
Sum squared deviation for Dave: 10
Sum squared deviation for Ivan: 44
Then divide by the number of data points to get the variance:
Variance for Mike: 12/5 = 3.2
Variance for Dave: 10/5 = 2
Variance for Ivan: 44/5 = 8.8
So according to the variance, Dave is the most consistent player.
Figure 7: Problem Consistency and its solution.
A.9 Student State Tracing Evaluation
$\mu$-average precision $\mu$-average recall $\mu$-average F1
Full version 0.82 0.75 0.77
No Justification 0.71 0.72 0.70
Short Memory 0.64 0.75 0.68
Methodology
We test the procedure using two tutoring conversations of about ten dialog turns for each of the two test problems detailed in Appendix A.7. These conversations are manually annotated to obtain the reference data. As detailed in Appendix A.2, we use an LLM to perform this task in our implementation of StratL. This LLM is prompted with the full conversation history and is asked to output a justification for the labels it selected. We compare this implementation to a version not outputting justifications and a version that only keeps the three last utterances in the prompts’ history to increase the processing speed.
As a multi-label classification task, we compute for each label $i$ the precision, recall, and F1 score. We aggregate these metrics across labels using an average weighted by the relative cardinality of each label, known as the micro average ($\mu$-average). We report the results in Table 6.
Results
A.10 V4 LLM Intent Selection Prompt
@@
2The tutor selects the intents based on the Productive Failure teaching method. The goal is to let the student explore the space of solution by themself to generate as many representation and solution methods as possible, even if they are only partially correct.
@@ 17 G_GREETINGS : Say goodbye to the student
A.11 User interface

### v2

@@ B.2.2 Intent Dependent Steering
B.2.3 V4 LLM Intent Selection Prompt
@@
2The tutor selects the intents based on the PF teaching method. The goal is to let the student explore the space of solution by themself to generate as many representation and solution methods as possible, even if they are only partially correct.
@@ 17 G_GREETINGS : Say goodbye to the student
B.3 Learning Sciences Rationales of our PF Modelling
The Intent Selection graph presented in Figure 2 has been created from expert knowledge in PF. We detail in this section some of the principles we used in its design.
In PF, the goal is to make the student explore the solution space of an open-ended problem as much as possible, even if they fail to produce a complete solution. This idea gave its name to Productive Failure: there is, in this case, a hidden efficacy to the student’s failure Kapur (2008). To this aim, in Figure 2 any mistake (coded by features ‘a’, ‘b’ or ‘c’) leads to a ‘Guide Self Correction’ intent. The aspiration is to make the student spot and correct their mistake by themself. If the error persists, we use the intent ‘Correct the Mistake’ that gives a strong hint to the student about their mistake, but without directly correcting it. If the mistake is computational only, we use the intent ‘Perform a Computation’. This allows the student to focus on the reasoning rather than spending time on computations. For the same reason, if the student asks for a basic theorem or definition they forgot (feature ‘h’), we use the intent ‘State’ to answer their question.
A good tutor should intervene very indirectly Merrill et al. (1992): a good principle to follow as a tutor is to give as little scaffolding as one can get away with. The intent used whenever a student manages to correct their mistake (arrow labeled ‘f’ leaving the ‘Guide Self-Correction’ node in Figure 2) or when the student suggests an intuitive or incomplete solution (state feature ‘d’) is the ‘Ask for the Next Step’ intent. This encourages the student to pursue their line of thought without heavier scaffolding.
One of the challenges of PF is to create a classroom culture that allows for failure and promotes psychological safety. Indeed, as students deal with open-ended problems before being given the tools to solve them, the experience can be stressful. In traditional classroom settings, the focus is directed towards getting the right solution and students are often evaluated on this with a grade. One of the most crucial design features of a PF problem session is to shift this goal from solving the problem to generating multiple RSMs regardless of their correctness Kapur and Bielaczyc (2012). As explained in Kapur and Bielaczyc (2012), “students are used to seeking assistance from their teachers so much so that they do so even before sufficiently trying to solve problems themselves. At the same time, teachers are just as used to providing assistance when it is asked for so much so that often opportunities for students to generate and explore RSMs are missed”. For this reason, in our modeling, if the student explicitly asks for help the intent ‘Ask for Intuition’ is triggered. This intent encourages the student to start with a guess or an intuition with no fear of being incorrect. Once the student is successful in providing an intuitive solution, or if they confidently propose an incomplete one (feature ‘e’), then the ‘Ask for Articulation’ intent is used to encourage an RSM creation by formulating their idea in concrete math. To mitigate stress, that can hinder the student’s learning Lepper et al. (1993), we trigger a ‘Give a Hint’ intent to ease the student’s progress if they show signs of lack of motivation or self-confidence.
A good PF problem allows for multiple solutions (and often a preferred one, called canonical solution). Then, if the student manages to propose a non-canonical solution to the problem (feature ‘j’), they should be encouraged to continue exploring. The ‘Identify Limits’ intent is triggered to ask the student to identify the limitations of their solution, encouraging them to develop a new one.
If the student reaches the canonical solution (feature ‘k’), we use the intent ‘Guide Self-Reflection’. It has been shown that it is beneficial for learning to engage in self-reflection (i.e. looking back critically at the work done) after a problem session Katz et al. (2000). Following this turn, the tutor is made ready to end the conversation through the ‘Say Goodbye’ intent.